OpenStudy (anonymous):

The official width-to-length ratio of the United States flag is 1:1.9. If a United States flag was 9.5 feet long, how wide would it be?

OpenStudy (anonymous):

let w denote the actual width. we can rewrite 1:1.9 as 1/1.9=10/19 so, 10/19=w/9.5 or we can also write it as, 10/19=10w/95 try solve the equation to get the value of w.

OpenStudy (anonymous):

the width is 5?

OpenStudy (anonymous):

good.thats right.

OpenStudy (anonymous):

ok! thank you once again!

OpenStudy (anonymous):

no problem Egirl.